OpenStudy (chillhill):

Solve for x: 4^x - 6(2^x) + 8 = 0

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 4^x - 6(2^x) + 8 = 0 }\) \(\large\color{black}{ \displaystyle (2\cdot 2)^x - 6(2^x) + 8 = 0 }\) \(\large\color{black}{ \displaystyle 2^x\cdot 2^x - 6(2^x) + 8 = 0 }\) \(\large\color{black}{ \displaystyle (2^x)^2 - 6(2^x) + 8 = 0 }\)

OpenStudy (solomonzelman):

Then you can do the following substitution: \(\large\color{black}{ \displaystyle w=2^x }\)

OpenStudy (solomonzelman):

You will obtain a new equation: \(\large\color{black}{ \displaystyle w^2 - 6w + 8 = 0 }\) and this you can solve for w, and afterward substitute these solutions for w into \(w=2^x\) (note that any negative solution for w will be excluded because \(2^x\) is never negative)